MadamMystery:

Continue the series. 1 3 5 11 19 Don't type anything useless :(

Bvellla:

Dm me

Bvellla:

Ok whats 19+10?

Bvellla:

11 + 8 = ? 19 + 10 = ? 29 + 12 = ? 41 + 14 = ?

Bvellla:

You should get your answer

MadamMystery:

Thanks 😊

MadamMystery:

@bvellla wrote:
You should get your answer
Can you explain this? If you don't mind 🙂